OpenStudy (anonymous):

If sin(x) = 1/2, what is cos(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (anonymous):

first of all you need to find angles such that sin(x)=1/2

OpenStudy (anonymous):

How do I do that exactly? Sorry I am just terrible at math :/

OpenStudy (anonymous):

there are two such angles:  the first one is well-known - x=30 degrees the another is easy to find - 180deg - 30deg = 150 deg

OpenStudy (anonymous):

in both cases cos(x) and tan(x) are different. when x=30: \[\cos(x)=\sqrt3/2\] \[\tan(x)=\sqrt3/3\]

OpenStudy (anonymous):

when x=150 deg \[\cos(x)=-\sqrt3/2\] \[\tan(x)=-\sqrt3/3\]